# Clone Deterministic 3D Worlds with Geometrically-Regularized World Models

### v1

@@ Abstract
A world model is an internal model that simulates how the world evolves. Given past observations and actions, it predicts the future of both the embodied agent and its environment. Accurate world models are essential for enabling agents to think, plan, and reason effectively in complex, dynamic settings. Despite rapid progress, current world models remain brittle and degrade over long horizons. We argue that a central cause is representation quality: exteroceptive inputs (e.g., images) are high-dimensional, and lossy or entangled latents make dynamics learning unnecessarily hard. We therefore ask whether improving representation learning alone can substantially improve world-model performance. In this work, we take a step toward building a truly accurate world model by addressing a fundamental yet open problem: constructing a model that can fully clone and overfit to a deterministic 3D world. We propose Geometrically-Regularized World Models (GRWM), which enforces that consecutive points along a natural sensory trajectory remain close in latent representation space. This approach yields significantly improved latent representations that align closely with the true topology of the environment. GRWM is plug-and-play, requires only minimal architectural modification, scales with trajectory length, and is compatible with diverse latent generative backbones. Across deterministic 3D settings and long-horizon prediction tasks, GRWM significantly increases rollout fidelity and stability. Analyses show that its benefits stem from learning a latent manifold with superior geometric structure. These findings support a clear takeaway: improving representation learning is a direct and useful path to robust world models, delivering reliable long-horizon predictions without enlarging the dynamics module.
@@ 1 Introduction
Creating a high-fidelity, interactive clone of an environment purely from observational data has long been a central ambition in artificial intelligence. Such a clone can serve as a simulator for reinforcement learning agents (Hafner et al., 2020; Hao et al., 2025), enable task planning in robotics (Mendonca et al., 2023), and facilitate controllable content generation in games (Alonso et al., 2024; Quevedo et al., 2024). World models are the primary tools for achieving this goal: they aim to capture an environment’s dynamics, predict its future states, and simulate its evolution (Sutton, 1991; Ha & Schmidhuber, 2018; Schrittwieser et al., 2020; LeCun, 2022).
Most current world models focus on open-world settings (Quevedo et al., 2024; He et al., 2025; Ball et al., 2025), generating unconstrained and often random environments in which each simulation produces a different world. This approach is intended to enhance generalization by exposing agents to diverse training scenarios. However, the emphasis on randomness can lead to unstable dynamics, making such models ill-suited for applications that demand reliable prediction and precise planning in fixed tasks. In these cases, they often produce futures that are merely plausible, not faithful.
@@
  • Representation learning. Exteroceptive sensory data, such as images, are high-dimensional and encode complex, nonlinear mappings from underlying physical processes. This makes accurate future-state prediction difficult even under full observability. For example, a simple spatial translation in the physical world can correspond to a highly nonlinear trajectory in pixel space. Addressing this requires a representation space that faithfully encodes the underlying physical states while minimizing information loss and noise, thereby simplifying the subsequent dynamics modeling task.
  • Dynamics modeling. Even with an optimal representation, the dynamics model must capture a broad range of transition patterns — including 3D transformations, logical rules, causal dependencies, and temporal memory. The challenge lies in building a unified mechanism that can accurately model all these regularities.
These challenges are inherently coupled: a poor representation forces the dynamics model to operate in a noisy, entangled latent space, increasing prediction complexity and reducing generalization; conversely, a well-structured representation is of limited utility if the dynamics model cannot capture the full spectrum of transition patterns. Progress toward generic neural world simulation therefore demands a co-design of representation learning and dynamics modeling, ensuring that the latent space is both physically meaningful and optimally aligned with the predictive capabilities of the dynamics model.
@@
Yet, we find that achieving accurate long-horizon cloning of even simple deterministic 3D environments remains an open challenge. Across all state-of-the-art baselines we evaluated, none were able to maintain fidelity over extended horizons: small prediction errors accumulate rapidly, causing trajectories to diverge from reality after only a few steps. In contrast, when the dynamics model is provided with the environment’s underlying physical states — rather than high-dimensional exteroceptive signals such as images — it can produce remarkably accurate long-horizon predictions. This observation suggests that the effectiveness of a world model is fundamentally constrained by the structure of its latent representation space. This raises the central question: How can we build a self-supervised representation that aligns with the underlying physical states, enabling stable and accurate long-horizon predictions?
While most representation learning methods are trained on IID datasets, real-world robotic systems naturally acquire continuous sensory trajectories through interaction. These trajectories inherently encode geometric and temporal regularities that can be leveraged to improve representation quality(Földiák, 1991; Wiskott & Sejnowski, 2002; Goroshin et al., 2015; Chen et al., 2018). Recent work(Wang et al., 2024) has shown that applying geometric regularization to temporal trajectories can substantially improve the topological structure of a latent space for 3D objects, benefiting tasks such as semantic classification and pose estimation. We extend this idea to the domain of 3D environments, demonstrating that trajectory-based geometric regularization can significantly enhance latent representations and, in turn, markedly improve the effectiveness of world models.
This work aims to bridge the gap to oracle-level performance through unsupervised representation learning. We introduce Geometrically-Regularized World Models (GRWM) — a framework that learns a latent space mirroring the geometry of the true state manifold, without requiring access to the ground-truth states. At its core is a lightweight geometric regularization module that can be seamlessly integrated into standard autoencoders, reshaping their latent space to provide a stable foundation for effective dynamics modeling. By focusing on representation quality, GRWM offers a simple yet powerful pipeline for systematically improving world model fidelity.
@@
    We formalize the problem of high-fidelity cloning of deterministic environments, shifting the focus from open-world generation to reproducible fidelity, and curate dedicated datasets for this task.
@@
    We introduce Geometrically-Regularized World Models (GRWM), a general, unsupervised geometric regularization method that enhances representation quality. This approach validates the “representation matters” hypothesis by demonstrating that a well-structured latent space systematically improves the performance of various dynamics models without altering their architecture.
@@
    We demonstrate that by focusing on representation, GRWM achieves state-of-the-art performance in long-horizon trajectory prediction. Extensive qualitative and quantitative analyses validate its success stems from learning a superior geometric structure in the latent space.
@@ 2 Preliminary
The Next-State Prediction Problem. World models aim to perform next-state prediction: learning a function that predicts future observations given the current state and an action. An environment evolves through latent states $s_{t}$, which yield observations $o_{t}$ after actions $a_{t}$. In the partially observable case, $o_{t}$ gives only partial information about $s_{t}$, requiring integration of past history to form a belief state. In the fully observable case, $s_{t}\approx o_{t}$, yet predicting in observation space remains difficult because simple physical transitions (e.g., translation) correspond to complex, non-linear changes in high-dimensional pixel space. This necessitates a learned representation space that can linearize these dynamics.
@@
Our objective is to learn a world model $M$ that can faithfully clone a deterministic environment from purely observational data. The model is trained on a dataset $\mathcal{D}$ consisting of trajectories $\tau=\{(o_{1},a_{1}),\dots,(o_{T},a_{T})\}$, where $o_{t}$ is an observation (image) and $a_{t}$ is an action at timestep $t$. The deterministic environment assumption states that, for any initial state and sequence of actions, there is exactly one resulting sequence of observations. The fidelity of the learned model is assessed through its ability to generate long-horizon rollouts. Specifically, given a starting observation $o_{1}$ and an action sequence $\{a_{t}\}_{t=1}^{T}$, the model produces a rollout $\{\hat{o}_{t}\}_{t=1}^{T}$. At each timestep $t$, we compute the frame-wise Mean Squared Error (MSE) between the generated and ground-truth observations: $\text{MSE}(t)\;=\;\|o_{t}-\hat{o}_{t}\|_{2}^{2}.$ This produces an error curve $\{\text{MSE}(t)\}_{t=1}^{T}$ that reveals how prediction error accumulates over time.
@@
Latent Generative Models. We adopt a latent generative framework consisting of two stages: (1) an autoencoder that learns a compressed latent representation (Kingma & Welling, 2013; Higgins et al., 2017), and (2) a generative model that captures the dynamics in that latent space (Ha & Schmidhuber, 2018; Hafner et al., 2019; Bruce et al., 2024). Our work focuses on the representation stage, as its quality critically determines the downstream rollout performance.
Contrastive Learning. Pure reconstruction objectives often yield degenerate representations that ignore temporal cues. Contrastive learning principles (Hadsell et al., 2006; van den Oord et al., 2018; Chen et al., 2018; 2020; He et al., 2020; Wang & Isola, 2020; Yeh et al., 2022; Garrido et al., 2022; Wang et al., 2024) suggest a remedy: enforce similarity for related samples and repulsion for unrelated ones. We extend this idea temporally, targeting the specific aliasing and stability challenges of world modeling.
@@ 3 Geometrically-Regularized World Models
3.1 Motivation: Representation Matters
Figure 1: Representation quality is the primary bottleneck for long-horizon prediction. Frame-wise MSE on the Maze 3x3 dataset. (Left) An oracle model using ground-truth states (black dotted) achieves near-zero error, establishing a performance upper bound. In contrast, a standard VAE-based world model (blue dashed) accumulates error rapidly. Our GRWM (green solid) significantly closes this gap by learning a more structurally aligned latent space (Right Bottom), while the VAE’s representation remains disorganized (Right Top). For further details, see Section 4.
To motivate our focus on representation learning, we first conduct a revealing experiment on the Maze 3x3 environment. We establish an “oracle” world model by training a world model directly on the ground-truth underlying states (i.e., the agent’s coordinates). As shown in Figure 1, this oracle, which bypasses the challenge of perception, predicts future states with near-perfect accuracy. This demonstrates that high-fidelity cloning is feasible if a perfect representation is available.
However, when the same dynamics model is trained on the latent space of a standard VAE, its performance collapses. This gap highlights that the primary bottleneck is not the dynamics model, but the representation itself. The VAE’s latent space, optimized solely for reconstruction, is not structurally aligned with the environment’s state manifold, leading to catastrophic errors. As a preview, Figure 1 also shows that a regularized VAE, trained with the geometric priors we introduce next, significantly closes this performance gap. This experiment solidifies our central thesis: high-fidelity cloning requires a latent space that is structurally aligned with the environment’s true underlying state manifold.
Building on the insight, we propose Geometrically-Regularized World Models, which consists of two key components: (1) an architecture that incorporates temporal context, and (2) a regularization loss that explicitly enforces structure in the latent space. We describe each component in the following subsections.
@@ 3.2 Temporal Contextualize Architecture
Learning underlying states from a single frame is difficult due to perceptual aliasing, where distinct states can yield nearly identical observations. For example, two different positions in a maze may appear visually indistinguishable. To resolve this ambiguity and capture the true dynamics of the environment, the model requires temporal context.
@@
VAE with Temporal Aggregation. Each frame is first encoded independently with a 2D CNN. The resulting frame-level features are then aggregated by a causal Transformer with a sliding temporal window, ensuring that $z_{t}$ only attends to a limited range of past frames up to and including time $t$. This windowed design captures short-term temporal context while maintaining causality.

### v2

@@ Abstract
A world model is an internal model that simulates how the world evolves. Given past observations and actions, it predicts the future physical state of both the embodied agent and its environment. Accurate world models are essential for enabling agents to think, plan, and reason effectively in complex, dynamic settings. However, existing world models often focus on random generation of open worlds, but neglect the need for high-fidelity modeling of deterministic scenarios (such as fixed-map mazes and static space robot navigation). In this work, we take a step toward building a truly accurate world model by addressing a fundamental yet open problem: constructing a model that can fully clone a deterministic 3D world. 1) Through diagnostic experiment, we quantitatively demonstrate that high-fidelity cloning is feasible and the primary bottleneck for long-horizon fidelity is the geometric structure of the latent representation, not the dynamics model itself. 2) Building on this insight, we show that applying temporal contrastive learning principle as a geometric regularization can effectively curate a latent space that better reflects the underlying physical state manifold, demonstrating that contrastive constraints can serve as a powerful inductive bias for stable world modeling; we call this approach Geometrically-Regularized World Models (GRWM). At its core is a lightweight geometric regularization module that can be seamlessly integrated into standard autoencoders, reshaping their latent space to provide a stable foundation for effective dynamics modeling. By focusing on representation quality, GRWM offers a simple yet powerful pipeline for improving world model fidelity.
@@ 1 Introduction
Creating a high-fidelity, interactive clone of an environment purely from observational data has long been a central ambition in artificial intelligence. Such a clone can serve as a simulator for reinforcement learning agents (hafner2020dreamer; hao2025neural; sutton), enable task planning in robotics (mendonca2023structured; unitree), and facilitate controllable content generation in games (alonso2024diffusionworldmodelingvisual; oasis2024; bruce2024genie). World models are the primary tools for achieving this goal: they aim to capture an environment’s dynamics, predict its future states, and simulate its evolution (dyna; ha2018world; schrittwieser2020mastering; lecun2022path).
Most current world models focus on open-world settings (oasis2024; he2025matrix; genie3; sora), generating unconstrained and often random environments in which each simulation produces a different world. This approach is intended to enhance generalization by exposing agents to diverse training scenarios (jang2025dreamgen; alonso2024diffusionworldmodelingvisual; hafner2020dreamer; hafner2019learning; bruce2024genie). However, the emphasis on randomness can lead to unstable dynamics, making such models ill-suited for applications that demand reliable prediction and precise planning in fixed tasks. In these cases, they often produce futures that are merely plausible, not faithful.
@@
  • Representation learning. Exteroceptive sensory data, such as images, are high-dimensional and encode complex, nonlinear mappings from underlying physical processes (battaglia2018relational; wang2020understanding; tschannen2018recent; arats2021improving; kang2024far). This makes accurate future-state prediction difficult even under full observability (ha2018world; hafner2019learning). For example, a simple spatial translation in the physical world can correspond to a highly nonlinear trajectory in pixel space (tancik2020fourier; nonlinear1; nonlinear2). Addressing this requires a representation space that faithfully encodes the underlying physical states while minimizing information loss and noise, thereby simplifying the subsequent dynamics modeling task (wang2022denoised; fang2024towards; seo2023masked; karlsson2023predictive; battaglia2018relational; zhou2024dinowmworldmodelspretrained).
  • Dynamics modeling. Even with an optimal representation, the dynamics model must capture a broad range of transition patterns — including 3D transformations, logical rules, causal dependencies, and temporal memory (dynamics1; dynamics2; dynamics3; dynamics4; hao2025neural; chen2024diffusion; huang2025selfforcing). The challenge lies in building a unified mechanism that can accurately model all these regularities (dynamics1; dynamics4; dynamics5; dynamics6).
These challenges are inherently coupled: a poor representation forces the dynamics model to operate in a noisy, entangled latent space, increasing prediction complexity and reducing generalization; conversely, a well-structured representation is of limited utility if the dynamics model cannot capture the full spectrum of transition patterns (representation1; representation2; tschannen2018recent; blau2019rethinking; kouzelis2025eq; stab). Progress toward generic neural world simulation therefore demands a co-design of representation learning and dynamics modeling, ensuring that the latent space is both physically meaningful and optimally aligned with the predictive capabilities of the dynamics model (ha2018world; seo2023masked; hafner2020dreamer; dynamics5; chen2024deep).
@@
Yet, we find that achieving accurate long-horizon cloning of even simple deterministic 3D environments remains an open challenge. Across all state-of-the-art baselines we evaluated, none were able to maintain fidelity over extended horizons: small prediction errors accumulate rapidly, causing trajectories to diverge from reality after only a few steps. In contrast, when the dynamics model is provided with the environment’s underlying physical states [^1] We refer to this version as the oracle model, since it has access to the true underlying states of the environment. — rather than high-dimensional exteroceptive signals such as images — it can produce remarkably accurate long-horizon predictions. This observation suggests that the effectiveness of a world model is fundamentally constrained by the structure of its latent representation space, and thus raises a natural question: How can we build a self-supervised representation that aligns with the underlying physical states, enabling stable and accurate long-horizon predictions?
A natural next step is to investigate how to bridge the gap to oracle-level performance through unsupervised representation learning. We show that applying the temporal contrastive learning principle (foldiak1991learning; wiskott2002slow; cltt3; cltt5; cltt7; klindt2020towards) as a geometric regularization can effectively curate a latent space that better reflects the underlying physical state manifold, demonstrating that contrastive constraints can serve as a powerful inductive bias for stable world modeling; we call this approach Geometrically-Regularized World Models (GRWM). At its core is a lightweight geometric regularization module that can be seamlessly integrated into standard autoencoders, reshaping their latent space to better align with the underlying physical states and provide a stable foundation for effective dynamics modeling. Through extensive experiments, we demonstrate that improved representation geometry directly translates into higher long-horizon fidelity across multiple world models, validating the “representation matters” hypothesis. GRWM acts as a plug-and-play component that systematically unlocks long-horizon fidelity for state-of-the-art world models.
@@
    Through diagnostic experiment, we quantitatively demonstrate that the primary bottleneck for long-horizon fidelity is the geometric structure of the latent representation, not the dynamics model itself. Our “oracle” model, which achieves near-perfect accuracy using ground-truth states, confirms that high-fidelity cloning is feasible and highlights that discovering these underlying states is a critical path toward perfect world modeling.
@@
    We demonstrate the use of temporal contrastive principles as a geometric regularizer for a world model’s autoencoder. Our approach, termed Geometrically-Regularized World Models (GRWM), leverages these principles to explicitly constrain the latent space to align with the environment’s underlying state manifold, showing that contrastive constraints can serve as a strong inductive bias for stable and faithful long-horizon world modeling.
@@
    We formalize the problem of high-fidelity cloning in deterministic environments, shifting the research focus from plausible open-world generation to reproducible fidelity. We also curate dedicated datasets for this task, providing a clearer benchmark for studying world model fidelity.
2 Related Work
Video World Models.
Video world models aim to predict future visual observations of an environment conditioned on agent actions, often using latent generative modeling as a central approach (oasis2024; he2025matrix; parkerholder2024genie2). Latent generative world models (seo2023masked; high; ha2018world) first encode observations into a latent space via an autoencoder and then model transitions in this space to predict future frames. This paradigm is widely applicable (oasis2024; parkerholder2024genie2; genie3; bruce2024genie; sora) and has inspired recent works exploring various enhancements, including interactive controllability (i1; oasis2024; a2), architectural innovations for causality (yin2025causvid; chen2024diffusion) and persistence (framepack; a1; far; framepackv1; song2025history), and improved physical accuracy (p1; p2; p3). Despite these advances, the predominant focus has been on improving dynamics modeling, e.g., designing more expressive latent transition models (t1; huang2025selfforcing; song2025history; a1) and adding causal and relational inductive biases (mendonca2023structured; hao2025neural; chen2024diffusion). The autoencoder’s role is largely engineering: to provide a compressive latent embedding for the dynamics model [^2] Commonly, the autoencoder is trained with a small KL weight and additional perceptual or adversarial losses (high)., with most work focusing on improving reconstruction quality rather than shaping the latent geometry for stable long-horizon prediction.
Moreover, most video world models target open-world or stochastic environments, emphasizing diversity and plausibility over faithful reproduction of a unique trajectory (oasis2024; he2025matrix; genie3; bruce2024genie; sora). In contrast, deterministic environments, where the goal is to precisely clone the true sequence of states, have received less attention. Our work highlights that in such settings, representation quality can be the primary bottleneck: even strong dynamics models fail if the latent space is poorly structured, motivating methods like GRWM that explicitly regularize latent geometry.
Contrastive Learning.
Pure reconstruction objectives often yield degenerate representations that ignore temporal cues. Contrastive learning principles (drlim; oord2018representation; chen2018sparse; chen2020simple; he2020momentum; wang2020understanding; yeh2022decoupled; garrido2022duality; wang2024pose) suggest a remedy: enforce similarity for related samples and repulsion for unrelated ones. Several methods extend this to temporal data, such as Contrastive Learning Through Time (CLTT) (cltt1; cltt2; cltt3; cltt4; cltt5; cltt6; cltt7; cltt8; wang2024pose), where contrastive objectives are used to capture temporal coherence in sequential observations.
GRWM builds directly on this foundation. While these methods successfully learn temporally coherent representations, they typically use contrastive objectives directly for representation learning. In contrast, we employ temporal contrastive learning principles as a geometric regularizer within a world modeling framework. This regularization explicitly reshapes the latent space of a world model’s autoencoder to align with the underlying physical state manifold. Through this formulation, we reveal that contrastive constraints can serve as a powerful inductive bias for improving world model fidelity, validating that representation quality is a central factor in achieving stable and accurate long-horizon prediction.
@@ 3 Preliminary
The Next-State Prediction Problem. World models aim to perform next-state prediction: learning a function that predicts future observations given the current state and an action (BrysonHo69; ha2018world; sutton). An environment evolves through latent states $s_{t}$, which yield observations $o_{t}$ after actions $a_{t}$.In the partially observable case, $o_{t}$ gives only partial information about $s_{t}$, requiring integration of past history to form a belief state. In the fully observable case, $s_{t}\approx o_{t}$, yet predicting in observation space remains difficult because simple physical transitions (e.g., translation) correspond to complex, non-linear changes in high-dimensional pixel space. This necessitates a learned representation space that can linearize these dynamics.
@@
Our objective is to learn a world model $M$ that can faithfully clone a deterministic environment from purely observational data. The model is trained on a dataset $\mathcal{D}$ consisting of trajectories $\tau=\{(o_{1},a_{1}),\dots,(o_{T},a_{T})\}$, where $o_{t}$ is an observation (image) and $a_{t}$ is an action at timestep $t$. The deterministic environment assumption states that, for any initial state and sequence of actions, there is exactly one resulting sequence of observations. The fidelity of the learned model is assessed through its ability to generate long-horizon rollouts. Specifically, given a starting observation $o_{1}$ and an action sequence $\{a_{t}\}_{t=1}^{T}$, the model produces a rollout $\{\hat{o}_{t}\}_{t=1}^{T}$. At each timestep $t$, we compute the frame-wise Mean Squared Error (MSE) [^3] We use frame-wise MSE to measure prediction error because it directly reflects pixel-wise differences and allows straightforward visualization of error accumulation over time. Since the images are normalized in [0,1], MSE can be easily converted to PSNR if needed. between the generated and ground-truth observations: $\text{MSE}(t)\;=\;\|o_{t}-\hat{o}_{t}\|_{2}^{2}.$ This produces an error curve $\{\text{MSE}(t)\}_{t=1}^{T}$ that reveals how prediction error accumulates over time.
Latent Generative Models.
@@
We adopt a latent generative framework (seo2023masked; high; ha2018world) consisting of two stages: (1) an autoencoder that learns a compressed latent representation (kingma2013auto; higgins2017beta; vqvae), and (2) a generative model that captures the dynamics in that latent space (ha2018world; hafner2019learning; bruce2024genie). Our work focuses on the representation stage, as its quality critically determines the downstream rollout performance.
4 Representation as the Bottleneck
Figure 1: Representation quality is the primary bottleneck for world model fidelity. Frame-wise MSE on the Maze 3x3 dataset. (Left) An oracle model using ground-truth states (black dotted) achieves near-zero error, establishing a performance upper bound. In contrast, a standard VAE-based world model (blue dashed) accumulates error rapidly. Our GRWM (green solid) significantly closes this gap by learning a more structurally aligned latent space (Right Bottom), while the VAE’s representation remains disorganized (Right Top). For further details, see Section 6.
We first conduct a foundational diagnostic experiment on the Maze $\times3$ environment to identify the source of error in high-fidelity cloning. We construct an “oracle” world model by training directly on the ground-truth underlying states, i.e., the agent’s coordinates $(x,y,\theta)$. This oracle, trained with full supervision, is designed to completely bypass the challenge of perception. It consists of three components: (1) An encoder that maps a history of observations to the previous state: $\hat{s}_{t-1}=f_{\text{enc}}(o_{t-k:t-1}),$ where $s_{t}=(x_{t},y_{t},\theta_{t})$. (2) A dynamics model that predicts the current state from the previous state and action: $\hat{s}_{t}=f_{\text{dyn}}(\hat{s}_{t-1},a_{t}).$ (3) A decoder that generates the current observation $\hat{o}_{t}$ from the current latent state: $\hat{o}_{t}=f_{\text{dec}}(\hat{s}_{t}).$
As shown in Figure 1, this oracle predicts future states with near-perfect accuracy. This result is a key finding: it quantitatively proves two critical points: (1) high-fidelity cloning is entirely feasible, and (2) the dynamics model itself is not the bottleneck. The challenge is perception. This oracle’s success establishes a clear performance upper bound and highlights that discovering these underlying states is a critical path toward perfect world modeling.
However, this oracle’s success contrasts with standard methods. When the exact same dynamics model is trained on the latent space of a standard VAE, its performance collapses. This massive performance gap highlights that the primary bottleneck is not the dynamics model, but the representation itself. The VAE’s latent space, optimized solely for reconstruction, is not structurally aligned with the environment’s state manifold, leading to catastrophic errors.
As a preview, Figure 1 also shows that a regularized VAE, trained with the geometric priors we introduce next, significantly closes this performance gap. This experiment solidifies our central thesis: representation quality is the primary bottleneck for world model fidelity.
@@ 5 Geometrically-Regularized World Models
We next explore how to reduce the gap to oracle-level performance through unsupervised representation learning. Our approach applies the temporal contrastive learning principle (cltt3; cltt5; cltt7; klindt2020towards) as a geometric regularizer on the autoencoder of a world model. This regularizer shapes the latent space so that its geometry better matches the underlying state manifold, providing a strong inductive bias for stable long-horizon prediction.
@@ 5.1 Temporal Contextualize Architecture
Learning underlying states from a single frame is difficult due to perceptual aliasing (alias1; alias2; alias3; alias4), where distinct states can yield nearly identical observations. For example, two different positions in a maze may appear visually indistinguishable. This also clarifies a key difference in the oracle experiment: the oracle uses coordinates, which encode the global state directly, whereas images provide only local information. A single image cannot specify the global position in environments with heavy aliasing. To recover the true state, the model must aggregate information over time, since a sequence of observations can form a unique signature of the underlying global state (cltt3; cltt5; cltt7).
@@
VAE with Temporal Aggregation.
Each frame is first encoded independently with a 2D CNN. The resulting frame-level features are then aggregated by a causal Transformer (vaswani2017attention) with a sliding temporal window, ensuring that $z_{t}$ only attends to a limited range of past frames up to and including time $t$. This windowed design captures short-term temporal context while maintaining causality.